OpenStudy (anonymous):

How wide a rectangular strip of land with length 3/4 mi and area 1/2 square mi.? Then find the perimeter of the rectangular strip of land. HELP.!!!

OpenStudy (anonymous):

HELP PLEASE

OpenStudy (anonymous):

length is \(\frac{3}{4}\) and area is \(\frac{1}{2}\) then since area is length times width the width \[\frac{3}{4}W=\frac{1}{2}\] so \[W=\frac{1}{2}\div \frac{3}{4}\]

OpenStudy (anonymous):

thanks so much.!v (:

OpenStudy (anonymous):

yw let me know what you get for the width

OpenStudy (anonymous):

k

OpenStudy (anonymous):

I got 2/3

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

so that's the answr right.?

OpenStudy (anonymous):

so the length is \(\frac{3}{4}\) and the width is \(\frac{2}{3}\) now you have to find the perimeter

OpenStudy (anonymous):

that is the width, yes, but that is not what the question asks for it is asking for the perimeter

OpenStudy (anonymous):

so how would I get to the perimeter

OpenStudy (anonymous):

\[2W+2L\]

OpenStudy (anonymous):

or if you like \[2(W+L)\]

OpenStudy (anonymous):

let me know what you get